A party balloon is inflated with helium at a temperature of 15 degrees, and has a volume of 3.5x10-2 m3. The pressure of the gas inside the ballon is 1.05x105 Pa.
calculate the number of miles inside the balloon
1
Expert's answer
2018-05-24T09:52:08-0400
As the ideal gas law says PV=nRT, where P – is a pressure, V – is a volume, n – is the number of moles of gas, T – is a temperature and R – is an ideal gas constant. So n=PV/RT, where temperature is in K. T=273+15=288K. n=(1.05×〖10〗^5×3.5×〖10〗^(-2))/(8.31×288)≈1.536 moles.